(II) Multiple continuous random variables:

8.2 Let X and Y have joint density fXY(x,y) = cx^2y for x and y in the triangle defined by 0 < x < 1, 0 < y < 1, 0 < x + y < 1

and fXY(x,y) = 0 elsewhere.

a. What is c?

b. What are the marginals fX(x) and fY(y)?

c. What are E[X], E[Y], Var[X] and Var[Y]?

d. What is E[XY]? Are X and Y independent?
Answer #1

yCx²dyda - 1x² / ydy dx = 1 Ifnyddyda = 1 (a) f(wy) = Ca?y, ocet, beye si penyayon - Sifacereyan = Sess Jy ay de -4 scf y jdVLX) = € (m?) - 4€ (X)] = 0.0285 - 60.05) TV (X) = 0.0260 E (Y) = 548464) dy = Sy.ayci-yg’dy = 2541-y?- 3y +342)dy = 2 8 (9²da 2O 4-0 00 431 5 6 7 ky) = $sa$ 4ony) dy dre - flug. Goly dyda =40***y2 ay) du 26 5 27 ( = -2 57361-3)*dx = 25*241-2232+3x